Shonna and Desi went shopping and spent $84 altogether. Shonna spent $15 less than twice as much as Desi. How much did Shonna spend?

Answer by tutorcecilia(2152)

Shonna and Desi went shopping and spent $84 altogether.
Shonna= (2)(Desi)-$15
Desi=x
.
2x-15+x=84
3x-15=84
3x=84+15
3x=99
x=33
.
Desi = $33
Shonna=(2)(33)-15=51
.
Checking: Desi+Shonna=84
33+51=84